A pressure gauge is sensitive to pressure variations that are close to ±1 cm or above. Calculate minimum wavelength at which a wave of 10 cm amplitude will be detectable in 1000 m of water.

Note, the pressure does not register smaller pressure variations
Answer #1

Given ampltude=10 cm=0.1 m and distance of water=1000 m

frequency of the wave=10 hertz

velocity=1000/1=1000 m/s It is the velocity per unit time

wavelength=velocity/frequency=1000/10=100 m